# ReservoirTTA: Prolonged Test-time Adaptation for Evolving and Recurring Domains

### v1

@@ 5.2 Analyses and Ablations
Component-wise Ablation Analysis and Runtime. Table 4 quantifies the impact of ReservoirTTA’s components on CIFAR-100-C (CSC, CDC) and ImageNet-C (CCC), along with runtime relative to EATA. Incorporating the Model Reservoir (MR) enables domain-specific adaptation, yielding a 1.9% gain at visit 20 in CSC. The Style Reservoir (SR) offers limited gains in CSC/CDC, likely because the frequent and abrupt domain resets lead to easily separable styles, making a single embedding sufficient for domain identification. In contrast, SR proves more effective in CCC, where gradual shifts benefit from accumulated style history, yielding a 1.2% improvement. Ensembling via EMR shows marginal benefit in CSC/CDC due to near one-hot assignments, but provides a 1.5% boost in CCC by leveraging multiple models. Overall, ReservoirTTA improves accuracy by 2.1% (CSC), 2.7% (CDC), and 1.5% (CCC), with only a 1.3$\times$ runtime increase over EATA.
Sensitivity Analysis of ReservoirTTA Hyperparameters. Figure 3 highlights the importance of setting $K^{\text{max}}$, with classification error stabilizing for $K^{\text{max}}\geq 16$ on CIFAR100-C—consistent with its domain structure. Overestimating the number of domains helps avoid premature merging and improves specialization. Additional ablations on CIFAR100-C and CCC (see Appendix E) show that ReservoirTTA is robust to key hyperparameters: performance remains stable across style reservoir sizes $M$, and variations in source sample count or quantile $q_{\text{th}}$ for threshold $\tau$ in Equation 7 affect error by less than 1%. As shown in Appendix I, larger batch sizes improve performance, but ReservoirTTA consistently outperforms baselines across settings. Notably, it requires less reliance on weight ensembling or Fisher regularization, indicating strong inherent variance control.
@@ 6 Conclusion
In this appendix, we provide further details on related work (Appendix A), our theoretical analysis (Appendix B), our algorithm (Appendix C), style features quality (Appendix D), domain identification via online clustering (Appendix E), model reservoir initialization (Appendix F), baselines (Appendix G), implementation details (Appendix H), additional ablation studies (Appendix I), as well as extra qualitative (Appendix J) and quantitative results (Appendix K). Note: The numbering of tables, figures, and equations in this appendix continues from the main document.
@@ B.1 TTA Variance Bound via Source Weighted Ensembling and Fisher Regularization
Conclusion. By mathematical induction, Equation 17 holds for all $t\geq 1$. ∎
@@
$\mathbb{E}[\nabla\mathcal{L}_{\text{TTA}}(\theta_{i},\mathbf{x}_{i})\nabla\mathcal{L}_{\text{TTA}}(\theta_{i},\mathbf{x}_{i})^{\top}]\approx\bar{V}.$
@@
From Equation 17, the update rule for $\theta_{t}$ is:
@@
Additionally, consider the limit $\alpha\to 1$. In this case, the denominator approaches 00 and we can expand the geometric sum:
@@
which matches the ensembling formulation in Equation 16. ∎
@@
The source-weighted ensembling update controls the variance of $\theta_{t}$ over time, as shown in Equation 18, but introduces bias toward the source model $\theta_{0}$. In the extreme case $\alpha=0$, no adaptation occurs, resulting in maximal bias. As $\alpha\to 1$, the method recovers standard SGD, minimizing bias but allowing unbounded variance. In practice, setting $\alpha$ close to 1 balances low variance with limited bias.
Proof of Theorem 1 (Bound on Divergence Probability). Let $\theta_{t}$ denote the model parameters at time $t$, and let $\theta^{*}_{\text{Task}}$ be the task-specific optimum. Suppose that $\mathbb{E}[\theta_{t}]\to\theta^{*}_{\text{Task}}$ and $\|\mathbb{E}[\theta_{t}]-\theta^{*}_{\text{Task}}\|<\|\theta_{0}-\theta^{*}_{\text{Task}}\|$. Then for any threshold $\beta>\|\theta_{0}-\theta^{*}_{\text{Task}}\|$, the probability of divergence from the stability region is bounded by:
@@
[…]
@@ B.2 Comparison of Single Model TTA and Model Reservoir TTA
In Figure 4, we compare the adaptation trajectories of a single model TTA approach and our proposed Model Reservoir TTA framework under recurring continual domain shifts. Consider an example with three domains, $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$, each associated with stability radii $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$, respectively. Let $\theta^{*}_{\text{Task}_{i}}=\arg\min_{\theta}\mathcal{L}_{\text{Task}_{i}}(\theta)$ for $i=1,2,3$ denote the task-optimal parameters that minimize the latent task loss in each domain. If the shift between optimal parameters exceeds the stability radius for a transition,
@@
then adaptation on $\mathcal{D}_{1}$ yields parameters near $\theta^{*}_{\text{Task}_{1}}$, but a subsequent shift to $\mathcal{D}_{2}$ may cause the adapted parameters to drift outside the stability region of $\mathcal{D}_{2}$. This drift is compounded in a recurring continual TTA setting, where the test stream follows $\mathcal{D}_{1}\to\mathcal{D}_{2}\to\mathcal{D}_{3}\looparrowright\mathcal{D}_{1}.$ The left panel of Figure 4 illustrates that even with sample filtering and weight ensembling, a single model TTA approach accumulates error at each domain transition and ultimately fails to re-adapt properly (i.e., the trajectory drifts away from the $\mathcal{D}_{1}$ optimum, converging instead near the source model). In contrast, the right panel demonstrates that our Model Reservoir TTA framework maintains separate, domain-specific trajectories that remain bounded within their respective stability regions, thereby enabling efficient re-adaptation when a previously encountered domain reoccurs.
@@
19:     Compute distance $\Delta=\min_{1\leq k\leq K_{t}}\|s_{t}-c^{k}\|$ See Equation 7
@@
26:     Compute soft assignment matrix $Q_{t}\in\mathbb{R}^{M\times K_{t}}$ See Equation 8
@@
28:        Update $c^{k}$ by gradient descent on $\mathcal{L}_{\mathrm{MI}}(Q_{t})$ See Equation 10
@@
Layer Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Adaptation Step $\xrightarrow{\hskip 22.76228pt}$
@@ Appendix C Algorithm of ReservoirTTA
For clarity, we detail the complete ReservoirTTA workflow in Algorithm 1.
@@ Appendix D Style Features Quality
Table 5 reports an ablation study on the impact of various VGG19 layer configurations for the extraction of style features. VGG19 is organized into four hierarchical blocks from which we extract style features at specific layers: Shallow ($[2,5,7]$), Middle-1 ($[10,12,14,16]$), Middle-2 ($[19,21,23,25]$), and Deep ($[28,30,32,34]$), where the indices denote the corresponding layers in the network. Additionally, we consider mixed configurations that integrate layers across different depths: Mixed-1 (Shallow + Middle-1), Mixed-2 (Deep + Middle-2), Mixed-3 (Shallow + Middle-1 + Middle-2 + Deep), and Mixed-4 (Shallow + Deep). We evaluated the four main configurations above (Shallow, Middle-1, Middle-2, Deep) along with several mixed configurations on three scenarios: recurring CSC and CDC on CIFAR100-C and CCC on ImageNet-C. The low standard deviations in the average error rates confirm that ReservoirTTA is robust to the choice of intermediate VGG19 layers.
@@
Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Adaptation Step $\xrightarrow{\hskip 22.76228pt}$
@@
Table 6 reports the ablation study on style feature representations in ReservoirTTA. The average classification error rates (%) for five different operations—mean, var, logvar, [mean, var], and gram—are evaluated. Notably, our logvar-based style representation delivers superior performance compared to the other operations.
Figure 5(a) presents the t-SNE visualization of style features extracted from three corruption benchmarks—CIFAR-10-C, CIFAR-100-C, and ImageNet-C—using our ReservoirTTA method. The datasets cover 15 distinct domains, which we organize into four categories: Noise (Gaussian Noise, Shot Noise, Impulse Noise), Blur (Defocus Blur, Glass Blur, Motion Blur, Zoom Blur), Weather (Snow, Frost, Fog, Brightness), and Digital (Contrast, Elastic, Pixelate, JPEG). As observed, samples from the same domain form well-defined clusters, while features from different domains remain clearly separated. This result confirms that style features effectively capture the inherent domain differences in these challenging corruption scenarios.
In Figure 5(b), we compare the feature distributions produced by our method, ReservoirTTA, against two baseline approaches, DPCore and CoLA, under the recurring CSC setting in the CIFAR-100-C dataset. We show the evolution of the t-SNE plots at the first, 10th, and 20th visits, illustrating how the features adapt as domain shifts accumulate over time. Our method consistently maintains well-separated and dense clusters throughout the adaptation process, while the baselines exhibit less distinct clustering. This comparison further demonstrates the effectiveness of ReservoirTTA in maintaining domain distinctions even in evolving environments.
@@ Appendix E Domain Identification via Online Clustering
Effect of Style Reservoir Capacity. The capacity of style reservoir $M$ determines how many style features are used to optimize the style centroids in Equation 9. As shown in Figure 6, in both recurring CSC and CDC settings, performance remains stable across different values of $M$, with low sensitivity to changes in this parameter. This suggests that even with an extremely small style reservoir size (e.g., $M=1$), our style-based domain assignment optimization remains robust and efficient, without incurring significant memory overhead. In CCC, increasing $M$ leads to a reduction in errors, confirming our ablation study analysis that a larger style reservoir allows for a more comprehensive observation of domain shifts, thereby improving performance. We set $M=1024$ for all main experiments.
@@
Effect of Number of Source Samples and Quantile Value. In Figure 7, we vary the number of unlabeled source samples from 300 to 10,000. For ROID*+ReservoirTTA, the error remains nearly constant, while for EATA+ReservoirTTA a slight decrease in error is observed with more source samples. In addition, minor fluctuations—especially in the recurring CSC and CDC settings. Figure 8 shows that varying the quantile value $q_{\text{th}}$ results in only minor error changes. A lower $q_{\text{th}}$ may spuriously trigger the detection of new domains, whereas a higher $q_{\text{th}}$ risks merging distinct domains. Overall, these results confirm that our style-based domain identifier is robust to variations in both the number of source samples and the quantile threshold, with only minimal impact on performance.
Ablation on Online Clustering Strategies. Table 7 compares different online clustering strategies (Sinkhorn–Knopp [1], Online K-Means, and our adaptive clustering scheme used for reservoirTTA) when integrated into ROID as the baseline. We evaluate recurring CSC and CDC on CIFAR-100-C, and CCC on ImageNet-C, reporting average classification error rates. Across all settings, our method consistently achieves the lowest error. In particular, for recurring CSC, our approach shows the best performance at visits 10 and 20, highlighting accurate domain partitioning and model assignment over multiple rounds. Sinkhorn–Knopp and Online K-Means provide moderate improvements but still trail behind our reservoir-based strategy. These findings confirm that our online clustering mechanism effectively separates target domains and assigns them to specialized models, leading to superior long-term adaptation.
Model Assignments per Domain. Figure 9 displays a heatmap illustrating how each target domain is assigned to different models in ReservoirTTA. The rows represent 15 corruption types (e.g., Gauss, Shot, Impulse, Defocus), while the columns correspond to 15 model indices. Each cell $[i,j]$ denotes the proportion of samples from the $i$-th domain that are routed to the $j$-th model, with row sums normalized to 1. We observe a clear one-to-one correspondence for most corruptions, though some exhibit similar style cues. Overall, the concentration along the diagonal reflects domain-specialized adaptation, whereas smaller off-diagonal entries highlight moderate sharing among visually related domains. This confirms that ReservoirTTA’s style-based clustering reliably assigns test samples to the most relevant domain-specific model.
Style Cluster Centroid Trajectories During Adaptation. Figure 10 shows the evolution (via trajectories) of style cluster centroids during adaptation. Each centroid is assigned a unique ID, which corresponds to the model index presented in Figure 9. As shown, each centroid remains closely linked to its assigned domain throughout adaptation, indicating that each model adapts exclusively to that domain.
FIFO vs. Reservoir Sampling for Domain Balancing. Figure 11 compares two sampling strategies—FIFO and Reservoir sampling—for maintaining a domain-balanced Style Reservoir in recurring CSC on CIFAR-100-C. It shows the distribution of different domains over time for two reservoir sizes ($M$=256 and $M$=1024). The results reveal that Reservoir sampling produces a more uniform and stable domain distribution than FIFO, even as the reservoir size varies. These observations underscore the robustness of our approach and support the use of Reservoir sampling as the preferred update strategy.
@@
Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Recurring visit $\xrightarrow{\hskip 22.76228pt}$ Adaptation Step $\xrightarrow{\hskip 22.76228pt}$
@@
Figure 12: Comparative study of model reservoir initializations in Recurring TTA. We compare two initialization strategies—one using source parameters and one based on mutual information (Equation 11)—across various datasets, TTA settings, and backbones. For CSC and CDC, error rates (%) are reported at the 20th recurrence; for CCC, we report the average error rate (%) over the full dataset.

### v3

@@ 5.2 Analyses and Ablations
Component-wise Ablation Analysis and Runtime. Table˜4 quantifies the impact of ReservoirTTA’s components on CIFAR-100-C (CSC, CDC) and ImageNet-C (CCC), along with runtime relative to EATA. Incorporating the Model Reservoir (MR) enables domain-specific adaptation, yielding a 1.9% gain at visit 20 in CSC. The Style Reservoir (SR) has limited effect on CSC/CDC—frequent, abrupt domain resets make styles easily separable, so a single embedding suffices—but helps in CCC, where gradual drift benefits from accumulated history (+1.2%). EMR ensembling is marginal on CSC/CDC (assignments are near one-hot) yet gives +1.5% on CCC by leveraging multiple models. As shown in Appendix˜I, reusing specialists via MR outperforms simple weight resets: RDumb’s blind reset 32.2% $\to$ domain-aware reset 31.2% $\to$ MR 28.4% (CSC, visit 20). Overall, ReservoirTTA improves accuracy by +2.1% (CSC), +2.7% (CDC), and +1.5% (CCC) at only $1.3\times$ EATA’s runtime.
Sensitivity Analysis of ReservoirTTA Hyperparameters. Figure˜3 highlights the importance of setting $K^{\text{max}}$, with classification error stabilizing for $K^{\text{max}}\geq 16$ on CIFAR-100-C consistent with its domain structure. Overestimating the number of domains helps avoid premature merging and improves specialization. Moreover, as shown in Figure˜4, ReservoirTTA is robust to batch-order variability, reliably estimating the number of domains after the first recurrence in the CSC setting even without constraining $K^{\text{max}}$. Across subsequent recurrences, the detected domain count increases marginally, consistent with the stable performance reported above. Additional ablations on CIFAR-100-C and CCC (see Appendix˜E) show that ReservoirTTA is robust to key hyperparameters: performance remains stable across style reservoir sizes $M$, and variations in source sample count or quantile $q_{\text{th}}$ for threshold $\tau$ in Equation˜7 affect error by less than 1%. As shown in Appendix˜I, larger batch sizes improve performance, but ReservoirTTA consistently outperforms baselines across settings. Notably, it requires less reliance on weight ensembling or Fisher regularization, indicating strong inherent variance control.
Additional Analysis. We further test two challenging TTA settings: (i) gradual shifts with smoothly varying severity, where ReservoirTTA adapts well to subtle transitions, and (ii) temporally correlated streams with category bias, where pairing with sampling/stabilization mitigates bias. Results (Appendix˜K) show effectiveness across diverse, realistic test-time scenarios.
Figure 4: Domain detection sensitivity of ReservoirTTA in recurring CSC on CIFAR-100-C. The plot shows the number of detected domains during the first recurrence (Left) and subsequent recurrences (Right) across different batch-order seeds. ReservoirTTA is largely insensitive to batch order, accurately estimating the number of domains, which remains stable over time.
@@ 6 Conclusion
A further limitation—common to most TTA methods—is updating parameters on every incoming batch regardless of convergence. Introducing an adaptive update trigger (to switch optimization on/off) could markedly reduce runtime overhead and improve practicality in resource-constrained settings.
@@
In this appendix, we provide further details on related work (Appendix˜A), our theoretical analysis (Appendix˜B), our algorithm (Appendix˜C), style features quality (Appendix˜D), domain identification via online clustering (Appendix˜E), model reservoir initialization (Appendix˜F), baselines (Appendix˜G), implementation details (Appendix˜H), additional ablation studies (Appendix˜I), as well as extra qualitative (Appendix˜J) and quantitative results (Appendix˜K). Note: The numbering of tables, figures, and equations in this appendix continues from the main document.
@@ B.1 TTA Variance Bound via Source Weighted Ensembling and Fisher Regularization
Conclusion. By mathematical induction, Equation˜17 holds for all $t\geq 1$. ∎
@@
$\mathbb{E}\Big[\mathrm{Var}[\nabla\mathcal{L}_{\text{TTA}}(\theta_{i},\mathbf{x}_{i})]\Big]\approx\bar{V}$
@@
From Equation˜17, the update rule for $\theta_{t}$ is:
@@
Additionally, consider the limit $\alpha\to 1$. In this case, the denominator approaches 0 and we can expand the geometric sum:
@@
which matches the ensembling formulation in Equation˜16. ∎
@@
The source-weighted ensembling update controls the variance of $\theta_{t}$ over time, as shown in Equation˜18, but introduces bias toward the source model $\theta_{0}$. In the extreme case $\alpha=0$, no adaptation occurs, resulting in maximal bias. As $\alpha\to 1$, the method recovers standard SGD, minimizing bias but allowing unbounded variance. In practice, setting $\alpha$ close to 1 balances low variance with limited bias.
Proof of Theorem˜1 (Bound on Divergence Probability). Let $\theta_{t}$ denote the model parameters at time $t$, and let $\theta^{*}_{\text{Task}}$ be the task-specific optimum. Suppose that $\mathbb{E}[\theta_{t}]\to\theta^{*}_{\text{Task}}$ and $\|\mathbb{E}[\theta_{t}]-\theta^{*}_{\text{Task}}\|<\|\theta_{0}-\theta^{*}_{\text{Task}}\|$. Then for any threshold $\beta>\|\theta_{0}-\theta^{*}_{\text{Task}}\|$, the probability of divergence from the stability region is bounded by:
@@
$\leq\|\theta_{t}-\mathbb{E}[\theta_{t}]\|+\|\mathbb{E}[\theta_{t}]-\theta^{*}_{\text{Task}}\|\quad\text{(triangle inequality)}.$ (23)
@@ B.2 Comparison of Single Model TTA and Model Reservoir TTA
In Figure˜5, we compare the adaptation trajectories of a single model TTA approach and our proposed Model Reservoir TTA framework under recurring continual domain shifts. Consider an example with three domains, $\mathcal{D}_{1}$, $\mathcal{D}_{2}$, and $\mathcal{D}_{3}$, each associated with stability radii $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$, respectively. Let $\theta^{*}_{\text{Task}_{i}}=\arg\min_{\theta}\mathcal{L}_{\text{Task}_{i}}(\theta)$ for $i=1,2,3$ denote the task-optimal parameters that minimize the latent task loss in each domain. If the shift between optimal parameters exceeds the stability radius for a transition,
@@
then adaptation on $\mathcal{D}_{1}$ yields parameters near $\theta^{*}_{\text{Task}_{1}}$, but a subsequent shift to $\mathcal{D}_{2}$ may cause the adapted parameters to drift outside the stability region of $\mathcal{D}_{2}$. This drift is compounded in a recurring continual TTA setting, where the test stream follows $\mathcal{D}_{1}\to\mathcal{D}_{2}\to\mathcal{D}_{3}\looparrowright\mathcal{D}_{1}.$ The left panel of Figure˜5 illustrates that even with sample filtering and weight ensembling, a single model TTA approach accumulates error at each domain transition and ultimately fails to re-adapt properly (i.e., the trajectory drifts away from the $\mathcal{D}_{1}$ optimum, converging instead near the source model). In contrast, the right panel demonstrates that our Model Reservoir TTA framework maintains separate, domain-specific trajectories that remain bounded within their respective stability regions, thereby enabling efficient re-adaptation when a previously encountered domain reoccurs.
@@
19:   Compute distance $\Delta=\min_{1\leq k\leq K_{t}}\|s_{t}-c^{k}\|$ See Equation˜7
@@
26:   Compute soft assignment matrix $Q_{t}\in\mathbb{R}^{M\times K_{t}}$ See Equation˜8
@@
28:    Update $c^{k}$ by gradient descent on $\mathcal{L}_{\mathrm{MI}}(Q_{t})$ See Equation˜10
@@
[…]
@@ Appendix C Algorithm of ReservoirTTA
For clarity, we detail the complete ReservoirTTA workflow in Algorithm˜1.
@@ Appendix D Style Features Quality
VGG-19 Configuration: Shallow vs. Deep. Table˜5 reports an ablation study on the impact of various VGG19 layer configurations for the extraction of style features. VGG19 is organized into four hierarchical blocks from which we extract style features at specific layers: Shallow ($[2,5,7]$), Middle-1 ($[10,12,14,16]$), Middle-2 ($[19,21,23,25]$), and Deep ($[28,30,32,34]$), where the indices denote the corresponding layers in the network. Additionally, we consider mixed configurations that integrate layers across different depths: Mixed-1 (Shallow + Middle-1), Mixed-2 (Deep + Middle-2), Mixed-3 (Shallow + Middle-1 + Middle-2 + Deep), and Mixed-4 (Shallow + Deep). We evaluated the four main configurations above (Shallow, Middle-1, Middle-2, Deep) along with several mixed configurations on three scenarios: recurring CSC and CDC on CIFAR100-C and CCC on ImageNet-C. The low standard deviations in the average error rates confirm that ReservoirTTA is robust to the choice of intermediate VGG19 layers. Table˜6 reports the ablation study on style feature representations in ReservoirTTA. The average classification error rates (%) for five different operations—mean, var, logvar, [mean, var], and gram—are evaluated. Notably, our logvar-based style representation delivers superior performance compared to the other operations.
@@
[…]
@@
Table 7: Effect of using the source backbone (w/o VGG) vs. frozen VGG-19 for style feature extraction on CIFAR-100-C (CSC) with ResNeXt-29. Error (%) across 20 recurrences.
Visits 1–10
Method 1 2 3 4 5 6 7 8 9 10
EATA 30.51 30.29 30.39 30.39 30.45 30.47 30.38 30.44 30.47 30.53
+ReservoirTTA w/o VGG 31.62 30.86 30.75 30.77 30.64 30.67 30.64 30.70 30.68 30.64
+ReservoirTTA 30.56 29.07 28.75 28.58 28.52 28.46 28.41 28.41 28.39 28.39
Visits 11–20
Method 11 12 13 14 15 16 17 18 19 20 Avg
EATA 30.51 30.46 30.47 30.51 30.51 30.48 30.51 30.54 30.47 30.47 30.46
+ReservoirTTA w/o VGG 30.65 30.65 30.61 30.60 30.61 30.61 30.56 30.53 30.52 30.51 30.69
+ReservoirTTA 28.38 28.40 28.37 28.37 28.37 28.40 28.36 28.43 28.37 28.38 28.57
Style Extractor Choice: VGG-19 vs. Source. We investigate whether using the source model for style feature extraction could replace the ImageNet-trained VGG. Following this suggestion, we replaced VGG with early layers of the source backbone for style calculation. On CIFAR100-C (CSC) with a ResNeXt-29 backbone, this swap increases mean error by +2.12% (Table˜7). Source backbones lack the broad texture priors learned from ImageNet and would make our approach architecture-specific. By contrast, VGG delivers consistent gains across ResNeXt-29, ViT-B-16, and ResNet-50 (GN) (see Appendix˜K). These results highlight that the VGG-19 style extractor with the automatically chosen threshold $\tau$ introduces minimal, dataset-agnostic overhead while providing substantially better domain detection and adaptation performance.
Style Features Quality. Figure˜6(a) presents the t-SNE visualization of style features extracted from three corruption benchmarks—CIFAR-10-C, CIFAR-100-C, and ImageNet-C—using our ReservoirTTA method. The datasets cover 15 distinct domains, which we organize into four categories: Noise (Gaussian Noise, Shot Noise, Impulse Noise), Blur (Defocus Blur, Glass Blur, Motion Blur, Zoom Blur), Weather (Snow, Frost, Fog, Brightness), and Digital (Contrast, Elastic, Pixelate, JPEG). As observed, samples from the same domain form well-defined clusters, while features from different domains remain clearly separated. This result confirms that style features effectively capture the inherent domain differences in these challenging corruption scenarios. In Figure˜6(b), we compare the feature distributions produced by our method, ReservoirTTA, against two baseline approaches, DPCore and CoLA, under the recurring CSC setting in the CIFAR-100-C dataset. We show the evolution of the t-SNE plots at the first, 10th, and 20th visits, illustrating how the features adapt as domain shifts accumulate over time. Our method consistently maintains well-separated and dense clusters throughout the adaptation process, while the baselines exhibit less distinct clustering. This comparison further demonstrates the effectiveness of ReservoirTTA in maintaining domain distinctions even in evolving environments.
@@ Appendix E Domain Identification via Online Clustering
Effect of Style Reservoir Capacity. The capacity of style reservoir $M$ determines how many style features are used to optimize the style centroids in Equation˜9. As shown in Figure˜7, in both recurring CSC and CDC settings, performance remains stable across different values of $M$, with low sensitivity to changes in this parameter. This suggests that even with an extremely small style reservoir size (e.g., $M=1$), our style-based domain assignment optimization remains robust and efficient, without incurring significant memory overhead. In CCC, increasing $M$ leads to a reduction in errors, confirming our ablation study analysis that a larger style reservoir allows for a more comprehensive observation of domain shifts, thereby improving performance. We set $M=1024$ for all main experiments.
@@
Effect of Number of Source Samples and Quantile Value. In Figure˜8, we vary the number of unlabeled source samples from 300 to 10,000. For ROID*+ReservoirTTA, the error remains nearly constant, while for EATA+ReservoirTTA a slight decrease in error is observed with more source samples. In addition, minor fluctuations—especially in the recurring CSC and CDC settings.
On the choice of the quantile threshold $q_{\text{th}}$: lowering $q_{\text{th}}$ increases sensitivity—detecting smaller shifts—but also raises false positives and spawns unnecessary domains; higher $q_{\text{th}}$ is more conservative and can delay detection of genuinely new domains. In a single sweep over the 15 CIFAR-100-C corruptions, $q_{\text{th}}=0.5$ severely overestimates (detects >100>100 domains), whereas $q_{\text{th}}=1.0$ merges similar corruptions—[zoom, motion, glass blur] and [Gaussian, shot, impulse noise]—yielding 9 domains. Intermediate thresholds balance sensitivity and precision: $q_{\text{th}}=0.99$ identifies 17 domains and $q_{\text{th}}=0.999$ identifies 14. Importantly, Figure˜9 shows that varying $q_{\text{th}}$ leads to only minor changes in error. Overall, our style-based domain identifier is robust to both the quantile setting and the number of source samples, with minimal impact on performance.
@@
[…]
@@
Table 9: Cosine vs. Euclidean vs. KL-divergence soft assignments for EATA+ReservoirTTA on CIFAR-100-C (CSC). Error (%) across 20 recurrences.
Visits 1–10
Variant 1 2 3 4 5 6 7 8 9 10
Cosine distance 31.94 31.04 30.91 30.81 30.76 30.77 30.74 30.77 30.60 30.58
Euclidean distance 31.76 30.37 30.19 29.96 29.81 29.93 29.92 29.84 29.89 29.86
KL divergence 32.01 30.59 30.14 30.16 30.19 30.00 30.04 30.09 29.99 30.05
Visits 11–20
Variant 11 12 13 14 15 16 17 18 19 20 Avg
Cosine distance 30.62 30.55 30.63 30.58 30.59 30.57 30.53 30.55 30.51 30.55 30.73
Euclidean distance 29.97 29.87 30.02 30.01 29.98 30.02 29.93 30.05 30.06 30.06 30.07
KL divergence 30.07 30.02 30.15 30.15 30.15 30.16 30.26 30.30 30.32 30.40 30.26
Ablation on Online Clustering Strategies. Table˜8 compares different online clustering strategies (Sinkhorn–Knopp [1], Online K-Means, and our adaptive clustering scheme used for reservoirTTA) when integrated into ROID as the baseline. We evaluate recurring CSC and CDC on CIFAR-100-C, and CCC on ImageNet-C, reporting average classification error rates. Across all settings, our method consistently achieves the lowest error. In particular, for recurring CSC, our approach shows the best performance at visits 10 and 20, highlighting accurate domain partitioning and model assignment over multiple rounds. Sinkhorn–Knopp and Online K-Means provide moderate improvements but still trail behind our reservoir-based strategy. These findings confirm that our online clustering mechanism effectively separates target domains and assigns them to specialized models, leading to superior long-term adaptation.
Ablation on Distance Metrics. We investigate alternative metrics for centroid assignment within EATA+ReservoirTTA. Euclidean distance aligns naturally with our quantile-based new-domain detector $\tau$, yet cosine similarity and KL divergence might, in principle, offer advantages in high-dimensional or streaming regimes. To assess this, we compare soft assignments using Euclidean distance, cosine similarity, and KL divergence on CIFAR-100-C (CSC) over 20 recurrences. As summarized in Table˜9, KL divergence outperforms cosine similarity—supporting its suitability for online adaptation—while Euclidean distance computed on log-variance features consistently achieves the lowest error across recurrences. These findings validate our choice to retain Euclidean distance and provide further empirical evidence of the method’s robustness, alongside a clear head-to-head comparison of distance metrics.
Model Assignments per Domain. Figure˜10 displays a heatmap illustrating how each target domain is assigned to different models in ReservoirTTA. The rows represent 15 corruption types (e.g., Gauss, Shot, Impulse, Defocus), while the columns correspond to 15 model indices. Each cell $[i,j]$ denotes the proportion of samples from the $i$-th domain that are routed to the $j$-th model, with row sums normalized to 1. We observe a clear one-to-one correspondence for most corruptions, though some exhibit similar style cues. Overall, the concentration along the diagonal reflects domain-specialized adaptation, whereas smaller off-diagonal entries highlight moderate sharing among visually related domains. This confirms that ReservoirTTA’s style-based clustering reliably assigns test samples to the most relevant domain-specific model.
Style Cluster Centroid Trajectories During Adaptation. Figure˜11 shows the evolution (via trajectories) of style cluster centroids during adaptation. Each centroid is assigned a unique ID, which corresponds to the model index presented in Figure˜10. As shown, each centroid remains closely linked to its assigned domain throughout adaptation, indicating that each model adapts exclusively to that domain.
FIFO vs. Reservoir Sampling for Domain Balancing. Figure˜12 compares two sampling strategies—FIFO and Reservoir sampling—for maintaining a domain-balanced Style Reservoir in recurring CSC on CIFAR-100-C. It shows the distribution of different domains over time for two reservoir sizes (MM=256 and MM=1024). The results reveal that Reservoir sampling produces a more uniform and stable domain distribution than FIFO, even as the reservoir size varies. These observations underscore the robustness of our approach and support the use of Reservoir sampling as the preferred update strategy.
Figure 13: Comparative study of model reservoir initializations in Recurring TTA. We compare two initialization strategies—one using source parameters and one based on mutual information (Equation˜11)—across various datasets, TTA settings, and backbones. For CSC and CDC, error rates (%) are reported at the 20th recurrence; for CCC, we report the average error rate (%) over the full dataset.